Which of the following gases under identical conditions (temperature, pressure) contains the 

largest number of molecules?

(A) 10 g nitrogen (B) 10 g chlorine (C) 10 g sulphur trioxide

(D) 10 g helium (E) 10 g carbon dioxide


1
Expert's answer
2021-05-10T01:25:48-0400

Under identical conditions (temperature, pressure) 1 mole of matter contains the equal number of molecules - 6.022×1023.

n = m / M

M (N2) = 28 g/mol

M (Cl2) = 70.9 g/mol

M (SO3) = 80.1 g/mol

M (He) = 4 g/mol

M (CO2) = 44 g/mol


n (N2) = 10 / 28 = 0.36 mol

n (Cl2) = 10 / 70.9 = 0.14 mol

n (SO3) = 10 / 80.1 = 0.13 mol

n (He) = 10 / 4 = 2.5 mol

n (CO2) = 10 / 44 = 0.23 mol

As 10 g of Helium contains the largest amount of moles - 2.5 moles, it has the largest number of molecules.